How do you solve the equation on the interval of [0,2pie) in radians? cos^2x+cos2x= 5/4
Answer by lwsshak3(11628)

How do you solve the equation on the interval of [0,2pie) in radians? cos^2x+cos2x= 5/4
..
cos^2x+cos2x= 5/4
2cos^2x=5/4
cos^2x=5/8
cosx=sqrt(5/8)=.79057
arccos(.79057)=.65906
x=.65906 and 5.62094 radians